Kamala got married 5 years ago.Today her age is 1 1/4 times her age at the time of marriage. Her sons age is1/10 times her present age. What is her son’s age?
Answer by Alex.33(110)

the current age of Kamala: y
1 1/4=5/4.
y/(5/4)=y-5
y=25
y/10=2.5
So her son's age is 2.5 years old (maybe 2 years old plus 6 months!)